for the following:
7*(2+3t)^2
I get the answer: 28 + 84t + 21t^2
what am I doing wrong??

Answer by lenny460(1073)

7(2 + 3t)^2
We use the order of operations - PEMDAS - Parentheses, Exponents, Multiplication, Division, Addition and Subtraction.
(2 + 3t)^2
We Multiply the above using the FOIL Method:
(2 + 3t)(2 + 3t)
2 * 2 = 4
2 * 3t = 6t
3t * 2 = 6t
3t * 3t = 9t^2
4 + 6t + 6t + 9t^2
Combine like terms:
6t + 6t = 12t
4 + 12t + 9t^2
7(4 + 12t + 9t^2)
7 * 4 = 28
7 * 12t = 84t
7 * 9t^2 = 63t^2
The Answer:
28 + 84t + 63t^2
or to put it in the correct order:
63t^2 + 84t + 28


obuong3@aol.com